Cost of Trade Credit What are the nominal and effective costs of trade credit under the credit terms of 1/20, net 35? Assume a 365-day year. Do not round intermediate calculations. Round your answers to two decimal places.

Nominal cost of trade credit: %

Effective cost of trade credit: %
Answer #1

Solution :-

Credit sterms are I, net 35 where, 1 represents discount percentage 20 represents 35 represents discount period days credit O365 35 -20 365 IS 99 - 0.010101010 x 24.33333333 = 0.2458 or 24.58% Effective cost of trade 365 days DCO-P credit Ell + DP DC= 1.2721 - 1 = 0.2771 or 27.71 % Thus, Nominal cost of trade Credit : 24.58% Effective cost of trade Credit: 27.71 %